OpenStudy (anonymous):

a METAL cube has a edge length of 11.4 mm and retriceof 667g. Calulate the density of metal...Please help I believe the answer is 148.cm^3....

OpenStudy (unklerhaukus):

the Volume of cube is \[V=s^3\] density is \[\rho=\frac mV\]

OpenStudy (anonymous):

yes but I need help find volume...ive gotten 1481544mm^3 but needs to convert to cm^3

OpenStudy (anonymous):

1481.544

OpenStudy (unklerhaukus):

to convert millimeters to centimeters \[1[\text{cm}]=10[\text{mm}]\] so \[1[\text{cm}]^3=(10[\text{mm}])^3=1000[\text{mm}]^3\]

OpenStudy (anonymous):

yes but I understand it I keep getten the answer wrong..can you please show me the procedure?

OpenStudy (anonymous):

the answer is 4.51g/cm^3 right??

OpenStudy (unklerhaukus):

\[\rho=\frac m{s^3}=\frac{667[\text g]}{{(11.4[\text{mm}])^3}}\]\[=\frac{667[\text g]}{1481.544[\text{mm}]^3}\]\[=\frac{667[\text g]}{1.481544[\text{cm}]^3}\]\[\approx 0.4502\tfrac{[\text g]}{[\text{cm}]^3}\]

OpenStudy (anonymous):

thanks